OpenStudy (anonymous):

Simplify (4 – 8i)(2 – 7i)

OpenStudy (anonymous):

4*2 +-7*4i-8*2i+56i^2

OpenStudy (anonymous):

this aint the answer

OpenStudy (anonymous):

??

OpenStudy (carson889):

Actually that is the answer. Just simplify: 4*2 = 8, -7i*4 = -28i, -8*2 = -16i and, -8i * -7i = 56i^2 = 56* (-1)^(1/2)^2 = -56 Therefore it equals: 8 - 28i - 16i - 56 = -44i - 48

OpenStudy (anonymous):

Umm, it looks like @carson889 answered it.